OpenStudy (anonymous):

for what value of x will the derivative of y = (sin2x)/4 + x/2 = 1? a. pi/3 b.O c.pi/2 d.pi/4

OpenStudy (anonymous):

i get answer 0 but can u check guys if its true

OpenStudy (anonymous):

is this \[y=\frac{\sin(2x)}{4}+\frac{x}{2}\]?

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

\[y'=\frac{\cos(2x)}{2}+\frac{1}{2}\]

sam (.sam.):

x=0

sam (.sam.):

cos(2x)=-1 2x=0 x=0

OpenStudy (anonymous):

then you want \[y'=\frac{\cos(2x)}{2}+\frac{1}{2}=0\] \[\cos(2x)=-1\] \[2x=\pi\] \[x=\frac{\pi}{2}\] is one answer

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

yes, it is C

OpenStudy (anonymous):

i feel not the answer is zero because a c and d have a pi then my prof is very tricky about question thx satelite

OpenStudy (anonymous):

@sam, cos(0)=1, not -1

sam (.sam.):

Im saying cos(2x)=-1 \[2x=\cos ^{-1}(-1)\] \[2x=\pi\] \[x=\pi/2\]

sam (.sam.):

@satellite73

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

y' = 1/2 (1 + cos(2x)) = 1; cos(2x) = 1; 2x = 0; x = 0°

OpenStudy (anonymous):

again !?

OpenStudy (nenadmatematika):

the answer is 0 I'm sure but be careful, don't write 0 degrees, it's 0 radians...so just write 0

OpenStudy (nenadmatematika):

just to be sure here the question is when will the derivative of the function y be equal to 1?